OpenStudy (j2lie):

How do you factor 7m^2+6m-1?

OpenStudy (unklerhaukus):

Use the quadratic formula\[ax^2+bx+c=0\]\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]\[(x-x_1)(x-x_2)=0\]

OpenStudy (j2lie):

What is the quadratic formula?

OpenStudy (unklerhaukus):

my second line is the quadratic formula

OpenStudy (j2lie):

isn't that for when you're trying to find the slope of a line.

OpenStudy (unklerhaukus):

The quadratic formula is used for finding the solutions \(x_1, x_2\) to the quadratic polynomial (a parabola), but you can also use if for factoring

OpenStudy (j2lie):

A parabola can be found using trinomial?

OpenStudy (unklerhaukus):

yes

OpenStudy (j2lie):

how do you do that?

OpenStudy (unklerhaukus):

the first step is to realise that \(m\) is like \(x\) now compare your expression , with my top line, what are \(a,b,c\)?

OpenStudy (j2lie):

Can you like give me an example using numbers?

OpenStudy (whpalmer4):

If you wanted to actually factor without solving first: \[7m^2+6m-1\]7*-1 = -7, find factors of -7 that sum to 6. -1,7 are said factors. \[7m^2 +7m -1m - 1\]\[(7m^2+7m)-(m+1)\]\[7m(m+1)-1(m+1)\]\[(m+1)(7m-1)\]

OpenStudy (j2lie):

oh I get it now. That's what I did and I got the same answer.

OpenStudy (unklerhaukus):

say you had this expression \(y^2-2y-3\) comparing with the polynomial we see that \(a=1\), \(b=-2\), \(c=-3\) now plugging these values into the quadratic formula \[y_{1,2}=\frac{-(-2)\pm\sqrt{(-2)^2-4(1)(-3)}}{2(1)}\\\qquad=\frac{2\pm\sqrt{4+7}}{2}\\\qquad=\frac{2\pm\sqrt{16}}{2}\\\qquad=\frac{2\pm4}{2}\\\qquad=\frac{2+4}2,\frac{2-4}{2}\\\qquad=\frac{6}{2},\frac{-2}{2}\\\qquad=3,-1\\~\\ y^2-2y-3=(y-3)(y+1)\]

OpenStudy (anonymous):

Another way to factor this is to notice that: \[7m^2 +6m-1 = 7m^2 +(7-1)m-1=(7m^2 +7m)-(m-1) \] =\[7m(m+1)-(m+1)=(7m-1)(m+1)\]

OpenStudy (whpalmer4):

So a parabola (which this is) will have crossings of the x-axis at the places where the product terms = 0, or \(m = -1, m = 1/7)\). Because the coefficient of the \(m^2\) term is positive, it is a parabola that opens upward. The vertex will be on a line perpendicular to the x-axis, midway between the two crossings of the x-axis, because parabolas are symmetric about the perpendicular line between vertex and directrix.

OpenStudy (anonymous):

Oops: The last (m-1) in the first line should be (m+1)

OpenStudy (whpalmer4):

changing (or forgetting to change) the sign in the second group is where I almost always make my mistakes in this, too :-)

OpenStudy (j2lie):

I forgot to multiply 7 and 1. And I thought it didn't make sense.

OpenStudy (whpalmer4):

Remembering the quadratic form as @UnkleRhaukus used is a good way to check your work, too. If you do a problem two different ways and get the same answer, the probability that you did it correctly is pretty good.

OpenStudy (j2lie):

I have never learned the first part of the quadratic formula.

OpenStudy (whpalmer4):

quadratic form is definitely the one I reach for if it's a real-world problem where the numbers aren't "nice" like they usually are in textbook examples.

OpenStudy (j2lie):

Yes the numbers aren't nice even using quadratic formula or the Magic X.

OpenStudy (unklerhaukus):

I always use the quadratic formula because it is one of the few formulae i have committed to memory and never have to look up to check, and it always works, Also i have trouble seeing how to split up that middle term, in the method used my whpalmer4 & RajeshRathod, it just isn't as intuitive (for me personally). But if you prefer that method and can see the factors as quickly as them, i would recommend you use that method. I would also note that if you use the quadratic formula method then there is sometimes another step (as in this case) you would get \(m_{1,2}=\dfrac{-6\pm8}{14}=\dfrac{2}{14},\dfrac{-14}{14}=\frac17,-1\) so the factor would come out as \[\qquad (m-\tfrac1{7})(m+1)=0\]multiplying by seven, to get rid of the fraction \[\qquad(7m-1)(m+1)=0\]

OpenStudy (j2lie):

ok.

OpenStudy (whpalmer4):

I do them just often enough to remember how, but not often enough to do them as quickly as I can factor things where the coefficient of the x^2 term = 1 :-) Being able to do it quickly depends on being able to factor the product of the leading and trailing coefficients in a hurry, of course. When the answer is an easily-recognizable prime number like it was for the original problem, that doesn't take long!

OpenStudy (j2lie):

ok.

OpenStudy (whpalmer4):

For \(y^2-2y-3\) we also have an easy case: product is -3, we need a sum to -2, so -3 and 1 are the choice. \[y^2-2y-3 = (y+1)(y-3)\]Or if you went to the trouble of splitting:\[y^2+y - 3y -3 =( y^2 + y )- (3y+3) = y(y+1)-3(y+1) = (y+1)(y-3)\]

OpenStudy (j2lie):

wow that looks complicated.

OpenStudy (whpalmer4):

it does? I thought that's what you've been doing all night!

OpenStudy (j2lie):

no I haven't learned that.

OpenStudy (whpalmer4):

Do a few hundred and it'll get easier :-)

OpenStudy (j2lie):

but I don't know how to do that.

OpenStudy (whpalmer4):

but that's what we did in the problem earlier...\(8x^2+32x-12x-48\) \[8x^2+32x -(12x+48) = 8x(x+4)-12(x+4) = (8x-12)(x+4)\]

OpenStudy (j2lie):

oh grouping it?

OpenStudy (whpalmer4):

Yes, that's what I was doing in the second line ("or if you went to the trouble of splitting")

OpenStudy (j2lie):

that's what you're suppose to do. you're suppose to split the middle coefficient.

OpenStudy (whpalmer4):

well, if you have to factor something where the coefficient of the \(x^2\) term is 1, you don't have to bother: you just factor the constant term, and pick the pair of factors that sums to the middle term. For example: \[x^2-6x+9\]Factors of 9 are 1*9,3*3,-1*-9,-3*-3 -3+-3 = -6, so that's the one we want. Factoring is\[x^2-6x+9 = (x-3)(x-3)\]

OpenStudy (whpalmer4):

Or \[x^2+5x+6\]Factors of 6 are 1*6, 2*3, 2+3 = 5, so 2 and 3 are what we want. \[x^2+5x+6 = (x+2)(x+3)\]

OpenStudy (whpalmer4):

As I was trying to illustrate before that other guy said I was doing it wrong, \[(ax+b)(x+c) = ax^2 + acx + bx + bc = ax^2 + (ac+b)x + bc\] We multiply first and last term to get \(abc\), then we look for factors of \(abc\) that give us \[ac+b = \text{<middle term>}\] For example: \[2x^2+5x+3\]Product of outer terms is 2*3 = 6. We need sum of factors to equal 5, so 2*3. We can't just write \[(x+2)(x+3)\]because the first term won't be \(2x^2\). So we need to write \[(2x+\text{<something>})(x+\text{<something else>})\]Because one of them gets multiplied by 2, we have to divide it by 2 before writing it in the factoring. We can't divide 3 by 2, so we divide 2 by 2 instead. A bit of trial-and-error here. \[(2x+3)(x+1) = 2x*x + 2x*1 + 3*x + 3*1 = 2x^2 + 5x + 3\]

OpenStudy (whpalmer4):

Pay attention to what happens as you multiply polynomials, and you'll have an easier time reversing the process.

OpenStudy (j2lie):

I get it now.

OpenStudy (whpalmer4):

Great!

OpenStudy (j2lie):

yay.